Edwin is four times as old as Joshua. In 12 years, Joshua's age will be half that of Edwin's. How old are they?
Answer by Maths68(1474)

Let
Present age of Joshua = j
Present age of Edwin = e
Edwin is four times as old as Joshua.
e=4j........(1)
In 12 years
e+12=j+12
then, Joshua's age will be half that of Edwin's
(e+12)*1/2=(j+12)
(e+12)=2(j+12)
e+12=2j+24
e=2j+24-12
e=2j+12
Put the value of e from (1) to above equation
e=2j+12
4j=2j+12
4j-2j=12
2j=12
2j/2=12/2
j=6
Put the value of j in (1)
e=4j
e=4*6
e=24
Present age of Joshua = j= 6 years old
Present age of Edwin = e = 24 years old